# SHyPar: A Spectral Coarsening Approach to Hypergraph Partitioning

### v1

@@ III-B1 Overview of Coarsening Refinement (HyperSF)
Figure 5 shows an overview of the HyperSF method . In this work, HyperSF is leveraged for only refining the most imbalanced node clusters (initially identified by HyperEF) with significantly smaller resistance diameters compared to the rest. Specifically, HyperSF aggregates strongly-coupled node clusters via minimizing Eq. (20): for each selected node set with large imbalance, HyperSF repeatedly solves a max $s$-$t$ flow, min $s$-$t$ cut problem to detect a set of neighboring node clusters that minimizes the local conductance HLC in Eq. (20). To this end, the following key steps are applied (as shown in Figure 6): (Step 1) an auxiliary hypergraph is constructed by introducing a source vertex $s$ and sink vertex $t$; (Step 2) each hyperedge is replaced with a directed graph; (Step 3) each seed node-set is iteratively updated by including new nodes into the set to minimize the HLC by repeatedly solving the max $s$-$t$ flow, min $s$-$t$ cut problem:
@@ III-B2 Local clustering algorithms
An algorithm is local if the input is a small portion of the original dataset. The HyperSF algorithm can be made strongly local when expanding the network around the seed nodes $\mathcal{C}$, which obviously benefits the proposed hypergraph coarsening framework: (1) applying the max $s$-$t$ flow, min $s$-$t$ cut problem on the local neighborhood of the seed nodes restricts node-aggregation locally and keeps the global hypergraph structure intact; (2) such a local clustering scheme significantly improves the algorithm efficiency due to the small-scale input dataset.
@@ III-B3 Flow-based Local Clustering in HyperSF
First, we apply HyperEF (Algorithm 2) to the hypergraph $H=(V,E,w)$ to compute a coarsened hypergraph $H^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ and identify isolated nodes, denoted by $\mathcal{C}$. To achieve flow-based local clustering of hypergraph nodes, HyperSF then constructs a sub-hypergraph $H^{\prime}_{L}$ by iteratively expanding the hypergraph around the seed node set $\mathcal{C}$ and then repeatedly solve the hypergraph cut problem to minimize HLC until no significant changes in local conductance are observed. Define $E^{\prime}(S)=\cup_{v^{\prime}\in V^{\prime},v^{\prime}\in\{e^{\prime}\}_{e^{\prime}\in E^{\prime}}}E^{\prime}(v)$ for any set $S\subseteq V^{\prime}$ and let $H^{\prime}_{\varsigma}=(V^{\prime}\cup\{s,t\},E^{\prime}\cup E^{\prime st})$, where $E^{\prime st}$ denotes the terminal edge set. HyperSF aims to construct a sub-hypergraph $H^{\prime}_{L}$ to replace $H^{\prime}_{\varsigma}$ that minimizes HLC by repeatedly solving a local version of (21). To this end, HyperSF sets up an oracle to discover a set of best neighborhood vertices for a given vertex $v^{\prime}$:
@@ III-D3 Hypergraph clustering with HyperSF
The flow-based clustering algorithm is a clustering refinement technique designed to enhance the clusters generated by HyperEF. It sorts the clusters based on their resistance diameters, starting with those that have low resistance diameters. HyperSF then creates new clusters by merging those with low-resistance diameters within the same neighborhood, utilizing a flow-based clustering technique. It is important to note that clusters with low resistance diameters typically indicate a strong connection among the nodes within the cluster, and therefore, merging these nodes is unlikely to significantly alter the global structure of the hypergraph. Ultimately, HyperSF aims to produce clusters characterized by low hypergraph local conductance (HLC) to preserve the spectral properties of the original hypergraph.

### v2

@@ III-B1 Overview of Coarsening Refinement (HyperSF)
Figure 5 shows an overview of the HyperSF method . In this work, HyperSF is leveraged for only refining the most imbalanced node clusters (initially identified by HyperEF) with significantly smaller resistance diameters compared to the rest. Utilizing the flow-based clustering method that is proposed in [67], HyperSF aggregates strongly-coupled node clusters via minimizing Eq. (20): for each selected node set with large imbalance, HyperSF repeatedly solves a max $s$-$t$ flow, min $s$-$t$ cut problem to detect a set of neighboring node clusters that minimizes the local conductance HLC in Eq. (20). To this end, the following key steps are applied (as shown in Figure 6): (Step 1) an auxiliary hypergraph is constructed by introducing a source vertex $s$ and sink vertex $t$; (Step 2) each hyperedge is replaced with a directed graph; (Step 3) each seed node-set is iteratively updated by including new nodes into the set to minimize the HLC by repeatedly solving the max $s$-$t$ flow, min $s$-$t$ cut problem:
@@ III-B2 Local clustering algorithms
The proposed flow-based clustering algorithm in [67] is strongly local by expanding the network around the seed nodes $\mathcal{C}$, which obviously benefits the coarsening framework: (1) applying the max $s$-$t$ flow, min $s$-$t$ cut problem on the local neighborhood of the seed nodes restricts node-aggregation locally and keeps the global hypergraph structure intact; (2) such a local clustering scheme significantly improves the algorithm efficiency due to the small-scale input dataset.
@@ III-B3 Flow-based Local Clustering in HyperSF
First, we apply HyperEF (Algorithm 2) to the hypergraph $H=(V,E,w)$ to compute a coarsened hypergraph $H^{\prime}=(V^{\prime},E^{\prime},w^{\prime})$ and identify isolated nodes, denoted by $\mathcal{C}$. Leveraging the introduced flow-based clustering in [67], HyperSF constructs a sub-hypergraph $H^{\prime}_{L}$ by iteratively expanding the hypergraph around the seed node set $\mathcal{C}$ and then repeatedly solve the hypergraph cut problem to minimize HLC until no significant changes in local conductance are observed. Define $E^{\prime}(S)=\cup_{v^{\prime}\in V^{\prime},v^{\prime}\in\{e^{\prime}\}_{e^{\prime}\in E^{\prime}}}E^{\prime}(v)$ for any set $S\subseteq V^{\prime}$ and let $H^{\prime}_{\varsigma}=(V^{\prime}\cup\{s,t\},E^{\prime}\cup E^{\prime st})$, where $E^{\prime st}$ denotes the terminal edge set. HyperSF aims to construct a sub-hypergraph $H^{\prime}_{L}$ to replace $H^{\prime}_{\varsigma}$ that minimizes HLC by repeatedly solving a local version of (21). To this end, HyperSF sets up an oracle to discover a set of best neighborhood vertices for a given vertex $v^{\prime}$:
@@ III-D3 Hypergraph clustering with HyperSF
The flow-based clustering algorithm introduced in [67] is used as a refinement technique to enhance the computed clusters by HyperEF. Sorting the clusters by resistance diameter, HyperSF begins with those that have the lowest values. It then applies a flow-based clustering technique to form new clusters by merging nearby clusters with low resistance diameters within the same neighborhood. It is important to note that clusters with low resistance diameters typically indicate a strong connection among the nodes within the cluster, and therefore, merging these nodes is unlikely to significantly alter the global structure of the hypergraph. Ultimately, HyperSF aims to produce clusters characterized by low hypergraph local conductance (HLC) to preserve the spectral properties of the original hypergraph.